# R-TPT: Improving Adversarial Robustness of Vision-Language Models through Test-Time Prompt Tuning

### v1

@@ 3.2 Refining Marginal Entropy Minimization
$minimize\ \mathcal{L}_{point}=\frac{1}{|\mathcal{B}|}\sum_{b=1}^{{|\mathcal{B}|}}\mathcal{H}({p^{b}}).$ (4)
@@ 4.1 Setup
CLIP-TeCoA [26] 79.3 44.3 66.9 15.8 10.2 1.0 30.8 9.0 6.6 6.6 24.5 10.7 14.5 10.8 34.6 6.7 33.4 13.1
Ensemble 72.7 55.1 59.9 38.9 5.6 2.7 26.6 16.0 4.2 4.2 23.5 16.2 12.5 11.0 26.4 14.0 28.9 19.8
TPT [38] 79.3 52.7 65.2 27.4 9.6 2.0 27.9 12.3 6.7 6.7 25.5 14.6 12.2 11.2 34.9 10.2 32.7 17.1
C-TPT [51] 79.8 47.3 66.1 19.5 10.6 1.3 29.4 10.7 6.4 6.4 26.2 12.4 13.0 11.1 36.4 8.1 33.5 14.6
MTA [54] 79.7 55.7 66.2 31.2 9.0 2.5 29.1 14.0 6.5 6.5 24.4 13.5 13.3 11.2 34.6 12.5 32.9 18.4
R-TPT 76.1 60.5 63.2 40.1 7.7 3.5 26.6 16.5 6.1 6.1 25.2 17.7 11.5 11.3 31.1 17.4 30.9 21.7
@@ 8 Experimental Results
CLIP [32] 91.4 0.2 85.1 0.0 60.1 0.0 64.0 0.0 18.1 18.1 43.0 0.0 35.8 0.0 61.6 0.0 57.4 2.3
TeCoA4 [26] 79.3 78.0 66.9 63.7 10.2 9.1 30.8 28.9 6.6 6.6 24.5 24.0 14.5 14.3 34.6 33.4 33.4 32.2
FARE4 [35] 86.3 85.4 76.7 73.8 39.2 34.4 37.0 34.0 9.5 9.5 28.3 27.3 16.6 16.3 44.2 41.9 42.2 40.3
APT4 [21] 10.7 0.4 10.0 0.2 1.5 0.1 0.9 0.2 2.6 2.6 9.0 0.1 7.8 6.7 3.7 0.2 5.8 1.3
APT4+TeCoA4 [21] 81.4 80.2 66.7 63.9 20.8 18.9 42.5 40.4 5.2 5.2 35.2 33.7 29.3 29.2 40.2 39.4 40.2 38.9
R-TPT 90.6 76.4 84.5 55.8 63.1 28.4 62.6 37.6 19.1 19.1 42.1 29.1 32.0 5.1 62.8 41.0 57.1 36.6

### v2

@@ 3.2 Refining Marginal Entropy Minimization
[…]
@@ 4.1 Setup
CLIP-TeCoA [26] 79.3 44.3 66.9 15.8 10.2 1.0 30.8 9.0 6.6 0.5 24.5 10.7 14.5 10.8 34.6 6.7 33.4 12.3
Ensemble 72.7 55.1 59.9 38.9 5.6 2.7 26.6 16.0 4.2 2.0 23.5 16.2 12.5 11.0 26.4 14.0 28.9 19.5
TPT [38] 79.3 52.7 65.2 27.4 9.6 2.0 27.9 12.3 6.7 1.7 25.5 14.6 12.2 11.2 34.9 10.2 32.7 16.5
C-TPT [51] 79.8 47.3 66.1 19.5 10.6 1.3 29.4 10.7 6.4 0.7 26.2 12.4 13.0 11.1 36.4 8.1 33.5 13.9
MTA [54] 79.7 55.7 66.2 31.2 9.0 2.5 29.1 14.0 6.5 1.6 24.4 13.5 13.3 11.2 34.6 12.5 32.9 17.8
R-TPT 76.1 60.5 63.2 40.1 7.7 3.5 26.6 16.5 6.1 2.7 25.2 17.7 11.5 11.3 31.1 17.4 30.9 21.2
@@ 8 Experimental Results
CLIP [32] 91.4 0.2 85.1 0.0 60.1 0.0 64.0 0.0 18.1 0.0 43.0 0.0 35.8 0.0 61.6 0.0 57.4 0.0
TeCoA4 [26] 79.3 78.0 66.9 63.7 10.2 9.1 30.8 28.9 6.6 5.9 24.5 24.0 14.5 14.3 34.6 33.4 33.4 32.2
FARE4 [35] 86.3 85.4 76.7 73.8 39.2 34.4 37.0 34.0 9.5 8.5 28.3 27.3 16.6 16.3 44.2 41.9 42.2 40.2
APT4 [21] 10.7 0.4 10.0 0.2 1.5 0.1 0.9 0.2 2.6 0.5 9.0 0.1 7.8 6.7 3.7 0.2 5.8 1.0
APT4+TeCoA4 [21] 81.4 80.2 66.7 63.9 20.8 18.9 42.5 40.4 5.2 5.0 35.2 33.7 29.3 29.2 40.2 39.4 40.2 38.8
R-TPT 90.6 76.4 84.5 55.8 63.1 28.4 62.6 37.6 19.1 9.2 42.1 29.1 32.0 5.1 62.8 41.0 57.1 35.3